OpenStudy (thatonegirl_):

The decibel level D of a sound is related to its intensity I by D=10log(I/I_0). If I_0 is 10^-12, then what is the intensity of a noise measured at 48 decibels? Express your answer in scientific notation, rounding to three significant digits, if necessary?

OpenStudy (danjs):

so solve \[48 = 10*\log(\frac{ I }{ 10^{-12} })\]

OpenStudy (thatonegirl_):

Yes

OpenStudy (danjs):

\[\frac{ 48 }{ 10 } = \log \frac{ I }{ 10^{-12} }\]

OpenStudy (danjs):

\[10^{\frac{ 48 }{ 10 }} = \frac{ I }{ 10^{-12} }\]

OpenStudy (thatonegirl_):

why does it turn to 10 when you divide by the log?

OpenStudy (danjs):

I = 10^(-12) * 10^(48/10)

OpenStudy (danjs):

no, recall ...

OpenStudy (danjs):

\[10^{\log(x)} = x\]

OpenStudy (thatonegirl_):

ohhh nvm sorry continue

OpenStudy (thatonegirl_):

u just rewrote it in exponential form, right?

OpenStudy (danjs):

to solve for I, you need to get rid of the log

OpenStudy (danjs):

by doing 10^(everything) on both sides

OpenStudy (danjs):

where 10 is the base of the log

OpenStudy (thatonegirl_):

okay i see

OpenStudy (danjs):

so you just have to calculate I = 10^(-12) * 10^(48/10)

OpenStudy (thatonegirl_):

so it comes out to be 6.31x10^-8 watt/m^2 ?

OpenStudy (danjs):

yeah, that is what i just got too

OpenStudy (danjs):

is that a reasonable range for the Intensity? I can't recall

OpenStudy (thatonegirl_):

okay cool, thanks a ton for all ur help!

OpenStudy (thatonegirl_):

idk lol its an answer choice. i plugged it back into the original equation and it works.

OpenStudy (danjs):

ha, k cool

OpenStudy (thatonegirl_):

thanks :D

OpenStudy (danjs):

np